# AMUN: Adversarial Machine UNlearning

### v1

@@ 7 Conclusions
Appendix A Related Works (cont.)
@@
Appendix B Implementation Details
@@
Appendix C AMUN + SalUn
@@
In our experiments, we try combining this idea with AMUN for updating a subset of the parameters that might be more relevant to the samples in $\mathcal{D_{\text{F}}}$. We refer to this combination as AMUNSalUn in Tables 1 and 2 in § 6.1 and Tables 7 and 8 in Appendix E. As the results show, AMUNSalUn constantly outperforms SalUn and for the cases that $\mathcal{D_{\text{R}}}$ is not available it also outperforms AMUN. In the setting where $\mathcal{D_{\text{R}}}$ is accessible, it performs comparable to AMUN. This is probably due to the fact that when $\mathcal{D_{\text{R}}}$ is not available and AMUN has access to only the samples $\mathcal{D_{\text{F}}}\cup\mathcal{D_{\text{A}}}$, SalUn acts as a regularization for not allowing all the parameters of the model that might not be relevant to $\mathcal{D_{\text{F}}}$ be updated. In the setting where $\mathcal{D_{\text{R}}}$ is available, involving it in fine-tuning will be a sufficient regularization that preserves models’ utility while unlearning $\mathcal{D_{\text{F}}}$.
@@
Appendix D Ablation Study (cont.)
In this section, we further discuss the ablation studies that were mentioned in § 6.2. We also present other ablation studies on using transferred adversarial examples (Appendix D.4) and weaker adversarial attacks (Appendix D.5) in AMUN.
D.1 Empirical Behavior of Retrained Models
@@
D.1.1 Confidence Values in Unlearned Models
@@
D.2 Adversarially Robust Models (cont.)
@@
D.3 Fine-tuning on Adversarial Examples (cont.)
As explained in § 6.2.1, we evaluate the effect of fine-tuning on test accuracy of a ResNet-18 model that is trained on CIFAR-10, when $\mathcal{D_{\text{A}}}$ is substituted with other datasets that vary in the choice of samples or their labels (see § 6.2.1 for details). In Figure 1 we presented the results when $\mathcal{D_{\text{F}}}$ contains $10\%$ of the samples in $\mathcal{D}$. We also present the results for the case where $\mathcal{D_{\text{F}}}$ contains $50\%$ of the samples in $\mathcal{D}$ in Figure 5. As the figure shows, even for the case where we fine-tune the trained models on only $\mathcal{D_{\text{A}}}$ which contains the adversarial examples corresponding to $50\%$ of the samples in $\mathcal{D}$ (right-most sub-figure), there is no significant loss in models’ accuracy. This is due to the fact that the samples in $\mathcal{D_{\text{A}}}$, in contrast to the other constructed datasets, belong to the natural distribution learned by the trained model. To generate the results in both Figures 1 and 5, we fine-tuned the trained ResNet-18 models on the all the datasets (see § 6.2.1 for details) for 20 epochs. We used a learning rate of $0.01$ with a scheduler that scales the learning rate by $0.1$ every 5 epochs.
@@
D.4 Transferred Adversarial Examples
@@ D.5 Weak Attacks
Appendix E Comparison Using Prior Evaluation Methods
@@
Appendix F Continuous Unlearning (cont.)

### v2

@@ 7 Conclusions
8 Acknowledgment
This work used Delta computing resources at National Center for Supercomputing Applications through allocation CIS240316 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program Boerner et al. (2023), which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Proofs
Here we provide the proof of Theorem 4.1:
Proof.
As we perform the unlearning by fine-tuning and performing a gradient descent update to $\theta_{o}$, we have: $\theta^{\prime}=\theta_{o}-\frac{1}{\beta}\nabla\hat{\mathcal{R}}(\theta_{o})$. Therefore, we can write:
$\|\theta^{\prime}-\theta_{u}\|_{2}^{2}$ $=\|\theta_{o}-\frac{1}{\beta}\nabla\hat{\mathcal{R}}(\theta_{o})-\theta_{u}\|_{2}^{2}$
$=\|\theta_{o}-\theta_{u}\|_{2}^{2}-\frac{2}{\beta}\langle\nabla\hat{\mathcal{R}}(\theta_{o}),\theta_{o}-\theta_{u}\rangle+\frac{1}{\beta^{2}}\|\nabla\hat{\mathcal{R}}(\theta_{o})\|_{2}^{2}$
$\leq\|\theta_{o}-\theta_{u}\|_{2}^{2}+\frac{2}{\beta}(\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta_{o}))+\frac{2}{\beta}(\hat{\mathcal{R}}(\theta_{o})-\hat{\mathcal{R}}(\theta^{\prime}))$
$=\|\theta_{o}-\theta_{u}\|_{2}^{2}+\frac{2}{\beta}(\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})),$
where the inequality is derived by using the smoothness property (^$\|\nabla\hat{\mathcal{R}}(\theta_{o})\|_{2}^{2}\leq 2\beta(\hat{\mathcal{R}}(\theta_{o})-\hat{\mathcal{R}}(\theta^{\prime}))$) and the convexity assumption which leads to the inequality: ^$\hat{\mathcal{R}}(\theta_{o}))\geq\hat{\mathcal{R}}(\theta_{u})+\langle\nabla\hat{\mathcal{R}}(\theta_{o}),\theta_{o}-\theta_{u}\rangle$.
Next, we derive an upper-bound for ^$\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})$ to replace in the above inequality. By the definition of unnormalized empirical loss on $\mathcal{D}^{\prime}$:
^$\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})$
$=\sum_{i=1}^{n-1}\ell(f_{\theta_{u}}(x_{i}),y_{i})+\ell(f_{\theta_{u}}(x),y)+\ell(f_{\theta_{u}}(x^{\prime}),y^{\prime})-\sum_{i=1}^{n-1}\ell(f_{\theta^{\prime}}(x_{i}),y_{i})-\ell(f_{\theta^{\prime}}(x),y)-\ell(f_{\theta^{\prime}}(x^{\prime}),y^{\prime})$
$=\ell(f_{\theta_{u}}(x),y)+\ell(f_{\theta_{u}}(x^{\prime}),y^{\prime})-\ell(f_{\theta^{\prime}}(x),y)-\ell(f_{\theta^{\prime}}(x^{\prime}),y^{\prime}),$
where the last equality was derived by the assumption that models are trained until they achieve near-00 loss on their corresponding dataset. Therefore, $\sum_{i=1}^{n-1}\ell(f_{\theta_{u}}(x_{i}),y_{i})=\sum_{i=1}^{n-1}\ell(f_{\theta^{\prime}}(x_{i}),y_{i})=0$ since the retrained model has been trained on the remaining samples and the unlearned model has been derived by a single step of gradient descent on the original model, that had been trained on $\mathcal{D}$.
To further simplify the derived terms above and reaching at our desired inequality, we focus on the term $-\ell(f_{\theta^{\prime}}(x),y)$. By adding and decreasing the term $\ell(f_{\theta_{o}}(x^{\prime}),y)$ we get:
$-\ell(f_{\theta^{\prime}}(x),y)$ $=-\ell(f_{\theta_{o}}(x^{\prime}),y)+\ell(f_{\theta_{o}}(x^{\prime}),y)-\ell(f_{\theta^{\prime}}(x),y)$
$\leq-\ell(f_{\theta_{o}}(x^{\prime}),y)+\ell(f_{\theta_{o}}(x^{\prime}),y)-\ell(f_{\theta_{o}}(x),y)-\langle\nabla\ell(f_{\theta_{o}}(x),y),\theta^{\prime}-\theta_{o}\rangle$
$=-\ell(f_{\theta_{o}}(x^{\prime}),y)+\ell(f_{\theta_{o}}(x^{\prime}),y)-\ell(f_{\theta_{o}}(x),y)$
$\leq-\ell(f_{\theta_{o}}(x^{\prime}),y)+L\delta,$
where the first inequality uses the convexity of the the loss function with respect to the parameters and the third derivations is due to the assumption that the original model achieves a zero loss on its training samples, including $(x,y)$ (hence, $\nabla\ell(f_{\theta_{o}}(x),y)=0$). The final inequality is due to the Lipschitzness assumption of model $f$ with respect to the inputs.
Appendix B Related Works (cont.)
@@ Appendix A Proofs
Appendix C Implementation Details
@@
Appendix D AMUN + SalUn
@@
In our experiments, we try combining this idea with AMUN for updating a subset of the parameters that might be more relevant to the samples in $\mathcal{D_{\text{F}}}$. We refer to this combination as AMUNSalUn in Tables 1 and 2 in § 6.1 and Tables 8 and 9 in Appendix F. As the results show, AMUNSalUn constantly outperforms SalUn and for the cases that $\mathcal{D_{\text{R}}}$ is not available it also outperforms AMUN. In the setting where $\mathcal{D_{\text{R}}}$ is accessible, it performs comparable to AMUN. This is probably due to the fact that when $\mathcal{D_{\text{R}}}$ is not available and AMUN has access to only the samples $\mathcal{D_{\text{F}}}\cup\mathcal{D_{\text{A}}}$, SalUn acts as a regularization for not allowing all the parameters of the model that might not be relevant to $\mathcal{D_{\text{F}}}$ be updated. In the setting where $\mathcal{D_{\text{R}}}$ is available, involving it in fine-tuning will be a sufficient regularization that preserves models’ utility while unlearning $\mathcal{D_{\text{F}}}$.
@@
Appendix E Ablation Study (cont.)
In this section, we further discuss the ablation studies that were mentioned in § 6.2. We also present other ablation studies on using transferred adversarial examples (Appendix E.4) and weaker adversarial attacks (Appendix E.5) in AMUN.
E.1 Empirical Behavior of Retrained Models
@@
E.1.1 Confidence Values in Unlearned Models
@@
E.2 Adversarially Robust Models (cont.)
@@
We also evaluated AMUN for unlearning in models that are adversarially trained. we performed our analysis on ResNet-18 models trained using TRADES loss Zhang et al. [2019] on CIFAR-10. We performed the experiments for unlearning 10% of the dataset in both cases where $\mathcal{D_{\text{R}}}$ is accessible and not. As the results in Table LABEL: show, in both settings AMUN is effective in unlearning the forget samples and achieving a low gap with the retrained models. This gap is obviously smaller when there is access to $\mathcal{D_{\text{R}}}$.
Unlearn Acc Retain Acc Test Acc FT AUC Avg. Gap
Retrain $82.33$$\pm 0.39$ $94.22$$\pm 0.21$ $81.72$$\pm 0.36$ $50.04$$\pm 0.34$ $0.00$
[…]
Table 5: Unlearning with access to $\mathcal{D_{\text{R}}}$. Evaluating AMUN when applied to ResNet-18 models trained using adversarial training. TRADES loss is used to train the models, and the unlearning is done on $10\%$ of CIFAR-10 Dataset ($\mathcal{D}$). Avg. Gap is used for evaluation (lower is better). The result has been reported in two cases: with and without access to $\mathcal{D_{\text{R}}}$. As the results show, AMUN is effective in both cases, with slight degradation in the more difficult setting of no access to $\mathcal{D_{\text{R}}}$.
E.3 Fine-tuning on Adversarial Examples (cont.)
As explained in § 6.2.1, we evaluate the effect of fine-tuning on test accuracy of a ResNet-18 model that is trained on CIFAR-10, when $\mathcal{D_{\text{A}}}$ is substituted with other datasets that vary in the choice of samples or their labels (see § 6.2.1 for details). In Figure 1 we presented the results when $\mathcal{D_{\text{F}}}$ contains $10\%$ of the samples in $\mathcal{D}$. We also present the results for the case where $\mathcal{D_{\text{F}}}$ contains $50\%$ of the samples in $\mathcal{D}$ in Figure 5. As the figure shows, even for the case where we fine-tune the trained models on only $\mathcal{D_{\text{A}}}$ which contains the adversarial examples corresponding to $50\%$ of the samples in $\mathcal{D}$ (right-most sub-figure), there is no significant loss in models’ accuracy. This is due to the fact that the samples in $\mathcal{D_{\text{A}}}$, in contrast to the other constructed datasets, belong to the natural distribution learned by the trained model. To generate the results in both Figures 1 and 5, we fine-tuned the trained ResNet-18 models on all the datasets (see § 6.2.1 for details) for 20 epochs. We used a learning rate of $0.01$ with a scheduler that scales the learning rate by $0.1$ every 5 epochs.
@@
E.4 Transferred Adversarial Examples
@@ E.5 Weak Attacks
Appendix F Comparison Using Prior Evaluation Methods
@@
Appendix G Continuous Unlearning (cont.)